A silver voltmeter,copper voltameter and oxygen voltameter are connected in series. Given that 4.32g of silver is deposited during the electrolytic process, determine(I) the mass of copper deposited(ii) the volume of oxygen liberated (Ag=108, Cu=64, O=16, G.M.V=22.4dm³




1
Expert's answer
2022-03-29T15:46:03-0400

Ag+ + Cu0 = Ag0 + Cu2+;

Ag+ +1e = Ag0 * 2

Cu0 -2e = Cu2+ * 1

Then,

2Ag+ +1e = 2Ag0;

Cu0 -2e = Cu2+;

And: 2Ag+ + Cu0 = 2Ag0 + Cu2+;

n(Ag) = m(Ag)/M(Ag) = 4.32/108 = 0.04 mol.

By the ionic reaction: n(Cu) = 1/2 * n(Ag) = 1/2 * 0.04 = 0.02 mol;

m(Cu) = n(Cu) * M(Cu) = 0.02 * 64 = 1.28 g.


Ag+ + O2- = Ag0 + O20;

Ag+ +1e = Ag0 * 4

2O2- -4e = O20 * 1

Then,

4Ag+ +1e = 4Ag0;

2O2- -4e = O20;

And: 4Ag+ + 2O2- = 4Ag0 + O20;

n(Ag) = m(Ag)/M(Ag) = 4.32/108 = 0.04 mol.

By the ionic reaction: n(O2) = 1/4 * n(Ag) = 1/4 * 0.04 = 0.01 mol;

V(O2) = n(O2) * Vm = 0.01 * 22.4 = 0.224 L.

Answer: m(Cu) = 1.28 g;

V(O2) = 0.224 L.